OpenStudy (anonymous):

Find the zeros of the function x^2 - x-4

OpenStudy (shadowfiend):

Do you know how to factor this equation?

OpenStudy (anonymous):

Yes but my problem is, is that I can't think of anything that added toger is 1 and multi to get 4

OpenStudy (shadowfiend):

Hm. Fair enough. So there are a couple of ways to approach this. Do you know the quadratic formula?

OpenStudy (anonymous):

yes -2 +- \[-b +_\sqrt{b^2 - 4ac} \over 2a\]

OpenStudy (shadowfiend):

Right. Do you know what a, b, and c are in your equation?

OpenStudy (anonymous):

yes a is 1x^2 b is 1x and c is 4

OpenStudy (shadowfiend):

Close, a, b, and c are the coefficients -- so a is 1, b is -1, and c is -4. Remember that a minus sign is part of the coefficient :) Can you plug those in?

OpenStudy (anonymous):

yeah I get \[1 \pm \sqrt{15} \over 2\]

OpenStudy (shadowfiend):

Right-o. And there are your two zeros.

OpenStudy (anonymous):

what are the two zeros?

OpenStudy (shadowfiend):

The quadratic equation describes two results: \[x = \frac{1 + \sqrt{15}}{2}\] and \[x = \frac{1 - \sqrt{15}}{2}\] Those are the two zeros of the equation.

OpenStudy (anonymous):

oh, ok thanks alot

OpenStudy (shadowfiend):

No problem, glad to help :)

OpenStudy (anonymous):

wouldn't this be 17 instead of 15?